A tank holds 95 liters of a chemical solution. Currently, the solution has a strength of 20%. How much of this should be drained and replaced with a 65% solution to get a final strength of 50%? Could you write the original equation also.
Answer by lwsshak3(11628)

A tank holds 95 liters of a chemical solution. Currently, the solution has a strength of 20%. How much of this should be drained and replaced with a 65% solution to get a final strength of 50%? Could you write the original equation also.
**
let x=amt of 20% solution to be drained and replaced with 65% solution.
95-x=amt of 20% solution remaining in tank.
..
20%(95-x)+65%x=50%*95
19-.2x+.65x=47.5
.45x=28.5
x=63+1/3 liters
amt of 20% solution to be drained and replaced with 65% solution=63+1/3 liters